Find the area of an equilateral triangle with side 12.
Answer by ewatrrr(24785)

 
Hi,

equilateral triangle with side of 12. | Note: Altitude bisects the base
h = sqrt(12^2 - 6^2) = sqrt(108) = 6sqrt(3)
A = (1/2)b*h = (1/2)12*6sqrt(3) = 36sqrt(3)